Solve the following system by substitution
5x-2y=-5
y-5x=3

Answer by ankor@dixie-net.com(22740)

Solve the following system by substitution
5x - 2y = -5
y - 5x = 3
:
since it has a 1y, use the 2nd equation for substituting;
y - 5x = 3
y = (5x+3)
:
Substitute (5x+3) for y in the 1st equation:
5x - 2y = -5
5x - 2(5x+3) = -5
:
Multiply what's inside the brackets:
5x - 10x - 6 = - 5
:
-5x = -5 + 6; added 6 to both sides
-5x = +1
x = 1/-5
x = -1/5
:
Remember we have y = 5x + 3; substitute 1/5 for x and find y:
y = 5(-1/5) + 3
y = -1 + 3
y = 2
:
:
Check our solutions in the 1st equation:
5x - 2y = -5
5(-1/5) - 2(2) =
-1 - 4 = -5